# Diffusion-Driven Semantic Communication for Generative Models with Bandwidth Constraints

### v1

@@ I Introduction
Deep learning-based semantic communication has recently demonstrated significant advantages for next-generation wireless communications, particularly in task-oriented communication systems designed to perform specific tasks or achieve particular objectives by extracting and transmitting only the relevant semantic information necessary. This approach has been applied to images [1, 2, 3], texts [4], and speech signals [5]. Neural networks can effectively extract useful information and handle complex data by learning the underlying features of the data. Numerous works have focused on improving communication efficiency. For example, deep joint source-channel coding (DJSCC)[6] merges source coding and channel coding into a single optimization problem with the goal of minimizing the overall transmission rate required to achieve a specific level of reliability. Compared to traditional wireless communication systems, semantic communication focuses on transmitting information in the semantic domain. Its primary goal is to ensure that the intended information of the message is accurately conveyed to the receiver.
Generative AI [7, 8] is becoming an essential application at the wireless edge. The generation models in generative AI have applications across various domains, including image, text, and audio generation. In image generation, these models can create new images based on given prompts or conditional inputs, effectively mapping latent features extracted by the model, to images. In wireless networks, image generation plays a crucial role in applications such as augmented reality (AR), virtual reality (VR), and mobile gaming through high-quality real-time visuals. Efficient image generation at the resource-limited devices reduces bandwidth usage and latency, vital for advanced applications like remote education and smart city surveillance. In semantic communication, the model leverages the advantages of semantic extraction and representation to transmit information at the semantic level. This process can also be seen as a conditional generation task under given channels and image features, i.e., generating corresponding images based on given conditions. Generative models over wireless networks benefit from semantic communication by focusing on transmitting essential features or information instead of raw data, which reduces data volume and enhances bandwidth efficiency. This approach also decreases latency, making it suitable for real-time applications, and improves robustness against noise and signal degradation.
@@
Recently, more work has focused on diffusion models for reducing the impact of wireless noise. The diffusion model [15, 16] adds Gaussian white noise during the forward process and gradually denoises the image during the backward process. This process can fit the communication system in an Additive White Gaussian Noise (AWGN) channel. For example, in [17] and [18], the diffusion model is employed to remove channel noise and improve reconstruction quality. The denoising diffusion probabilistic model (DDPM) is further applied in [19] for probabilistic constellation shaping in wireless communications. To address perception distortion in finite block lengths, diffusion is integrated into DJSCC in [20]. Moreover, some works use degraded signals based on diffusion models for semantic communication. A hybrid joint source-channel coding (JSCC) scheme is proposed in [21], where conventional digital communication with the compressed image is complemented with a generative refinement component to enhance the perceptual quality of reconstruction. Invertible Neural Networks are proposed in [22], where the signal is decomposed at the transmitter and is estimated from the degraded part using the diffusion models to recover high-quality source images under extreme conditions at the receiver. The diffusion models are used to denoise in wireless communication. However, current diffusion models overlook bandwidth constraints, focusing on denoising on the original image size or features in image processing and generation tasks.
@@
This paper introduces a diffusion-driven semantic communication framework with advanced VAE-based compression for bandwidth-constrained generative model. We design an architecture that utilizes the diffusion model, where the signal transmission process through the channel acts as the forward process in diffusion. To reduce bandwidth requirements, we incorporate a downsampling module and a paired VAE-based upsampling module with reparameterization at the receiver to ensure that the recovered features conform to the Gaussian distribution. We use the Stable Diffusion model to effectively enhance information transmission by integrating compression in the proposed bandwidth-constrained semantic communication. Our contributions can be summarized as follows:
@@
    We propose a communication-efficient generative semantic communication system, which incorporates the forward process of diffusion and channel noise, mapping the channel noise into the $T$-th forward processes in diffusion, adaptable to different signal to noise ratios (SNRs) of the wireless channel. At the receiver, we effectively utilize the diffusion model of the reverse process to remove the channel noise.
@@
    To reduce bandwidth requirements, we incorporate a downsampling module and a paired upsampling module to restore the signal at the receiver. Considering the potential distortion of Gaussian distribution, we introduce a VAE-based upsampling network with reparameterization at the receiver. Leveraging the reparameterization of the VAE ensures that the recovered features adhere to the Gaussian distribution necessary for the diffusion process.
@@ III-B Adaptive Forward Process
According to Eqn. (15), when the variance of noise in wireless channel is $\sigma^{2}$, the transmission process over the channel is equivalent to the $u$-th step forward process. This can be expressed as:
$\left.p_{c}\left(\hat{\boldsymbol{s}}\mid\boldsymbol{y}\right)\sim p\left(\boldsymbol{y}_{u}\mid\boldsymbol{y}_{0}\right)\right|_{\bar{\alpha}_{u}=1/\left(1+\sigma^{2}\right)}.$ (16)
@@
$\left.p_{d}\left(\hat{\boldsymbol{s}}_{u-1}\mid\hat{\boldsymbol{s}}_{u},\boldsymbol{\phi}\right)\sim p_{\boldsymbol{\phi}}\left(\boldsymbol{y}_{u-1}\mid\boldsymbol{y}_{u}\right)\right|_{\bar{\alpha}_{u}=1/\left(1+\sigma^{2}\right)},$ (17)
@@ IV The Details of the Diffusion-Driven Semantic Communication System with VAE-Based Compression
In this section, we present a diffusion-driven semantic communication framework enhanced by VAE-based compression. To ensure the Gaussian distribution of features, we employ VAE-based reparameterization for feature reconstruction. Additionally, we introduce guidance to focus on training the paired upsampling and downsampling networks to reduce training costs. The generator integrated with compression leverages guidance from the distribution of uncompressed generator, enabling the efficient generation of corresponding images with reduced bandwidth. Fig. 3 illustrates the framework for detailed semantic communication with compression and guidance.
@@ IV-C The Hybrid Loss from Reparameterization and Guidance
$\mathcal{L}_{v}=\lambda D_{\text{KL}}\left(q_{\boldsymbol{\psi},\boldsymbol{\omega}}\left(\hat{\boldsymbol{y}}\mid\!\boldsymbol{y}\right)\|p_{\boldsymbol{\phi}}\left(\hat{\boldsymbol{y}}\right)\right)+\mathbb{E}\!\left[\left({\boldsymbol{y}^{\prime}\!-\!\boldsymbol{y}}\right)^{2}\!\right],$ (22)
@@
Proof: The proof is along the lines of the proof in Appendix A and is therefore omitted..
@@ V Results
In this section, we present experimental results to validate the effectiveness of the proposed semantic communication generation with compression and guidance. The uncompressed generator employed in our approach is the stable-diffusion-v2 [16]. Our method is evaluated on the LSUN-Churches and LSUN-Bedrooms datasets [34]. We configure the number of diffusion timesteps to $T=1000$ with a linear schedule, set the learning rate to $1\times 10^{-4}$, and use the training batch size of 16. In addition, we also visualized the semantic distinctions between the generated images and real images. This helps to provide a more intuitive understanding of the effectiveness of generating semantic communication based diffusion. The simulations are conducted on the computer equipped with an Intel Xeon Silver 4110 CPU @ 2.10GHz and an NVIDIA RTX A40 GPU.
Figure 5: The visual results of the adaptation generation across various SNR levels. When the SNR is low, the semantic information of the image can be correctly transmitted, and as the SNR increases, more details can be recovered.
V-A The Metrics
@@ V-B Performance of Channel Integration and Adaptation
(a) PSNR of the proposed scheme.
(b) SSIM of the proposed scheme.
(c) LIPIPS of the proposed scheme.
(d) CLIP-score of the proposed scheme.
Figure 6: The comparisons of the proposed generation semantic communication on LSUN-Churches dataset across different SNRs.
@@
TABLE III: The adaptive generation in LSUN-Bedrooms datasets.
Metric 0dB 3dB 6dB 9dB 12dB
PSNR $\uparrow$ 22.67 24.16 25.51 26.70 27.73
SSIM $\uparrow$ 0.68 0.73 0.76 0.79 0.81
CLIP-score $\uparrow$ 0.88 0.90 0.91 0.91 0.91
FID $\downarrow$ 29.20 23.90 19.96 17.22 15.04
LPIPS $\downarrow$ 0.22 0.18 0.15 0.12 0.11
TABLE IV: The adaptive generation in LSUN-Churches datasets.
Metric 0dB 3dB 6dB 9dB 12dB
PSNR $\uparrow$ 20.10 21.33 22.46 23.46 24.31
SSIM $\uparrow$ 0.57 0.62 0.66 0.70 0.73
CLIP-score $\uparrow$ 0.86 0.88 0.89 0.89 0.90
FID $\downarrow$ 27.59 23.01 19.86 17.23 15.61
LPIPS $\downarrow$ 0.24 0.20 0.17 0.15 0.13
V-C Performance of Compression with Different Bandwidths
(a) PSNR of the proposed scheme.
(b) SSIM of the proposed scheme.
(c) LIPIPS of the proposed scheme.
(d) CLIP-score of the proposed scheme.
Figure 7: The Improvement of the reparameterization and guidance on LSUN-Churches dataset across different SNRs.
@@ V-C Performance of Compression with Different Bandwidths
In the uncompressed generator, the noise from the wireless channel is treated as a complete forward process to adapt to different SNR conditions. We also evaluate the generation performance on the LSUN Bedrooms and LSUN Churches datasets. The results are shown in Tables III and IV. The CLIP-score metrics, used for evaluating the semantic content of the images, demonstrate indicate that diffusion-driven semantic communication ensures accurate transmission of semantic information across various SNR levels. However, at higher SNR, diffusion-driven semantic communication not only ensures accurate transmission of semantic information but also better preserves the details of the images. This results is reflected in the pixel-level evaluation metrics such as PSNR and SSIM. The visual results of generated images at different SNR levels are illustrated in Fig. 5. Even at lower SNR, semantic information can be effectively transmitted. For instance, in the figure, at low SNR, there are differences in the details of the church’s roof, windows, and architectural style, but the overall semantic content remains consistent. With increasing SNR, these details are better recovered in the images.

### v2

@@ I Introduction
Deep learning-based semantic communication has recently demonstrated significant advantages for next-generation wireless communications, particularly in task-oriented communication systems designed to perform specific tasks or achieve particular objectives by extracting and transmitting only the relevant semantic information necessary. This approach has been applied to images [1, 2, 3], texts [4], and speech signals [5]. Neural networks can effectively extract useful information and handle complex data by learning the underlying features of the data. Numerous works have focused on improving communication efficiency. For example, deep joint source-channel coding (DJSCC)[6] merges source coding and channel coding into a single optimization problem with the goal of minimizing the overall transmission rate required to achieve a specific level of reliability. Compared to traditional wireless communication systems, semantic communication focuses on transmitting information in the semantic domain [7]. Its primary goal is to ensure that the intended information of the message is accurately conveyed to the receiver.
Generative AI [8, 9] is becoming an essential application at the wireless edge. The generation models in generative AI have applications across various domains, including image, text, and audio generation. In image generation, these models can create new images based on given prompts or conditional inputs, effectively mapping latent features extracted by the model, to images. In wireless networks, image generation plays a crucial role in applications such as augmented reality (AR), virtual reality (VR), and mobile gaming through high-quality real-time visuals. Efficient image generation at the resource-limited devices reduces bandwidth usage and latency, vital for advanced applications like remote education and smart city surveillance. In semantic communication, the model leverages the advantages of semantic extraction and representation to transmit information at the semantic level [10]. This process can also be seen as a conditional generation task under given channels and image features, i.e., generating corresponding images based on given conditions. Generative models over wireless networks benefit from semantic communication by focusing on transmitting essential features or information instead of raw data, which reduces data volume and enhances bandwidth efficiency. This approach also decreases latency, making it suitable for real-time applications, and improves robustness against noise and signal degradation.
@@
Recently, more work has focused on diffusion models for reducing the impact of wireless noise. The diffusion model [17, 18] adds Gaussian white noise during the forward process and gradually denoises the image during the backward process. This process can fit the communication system in an Additive White Gaussian Noise (AWGN) channel. For example, in [19] and [20], the diffusion model is employed to remove channel noise and improve reconstruction quality. The denoising diffusion probabilistic model (DDPM) is further applied in [21] for probabilistic constellation shaping in wireless communications. To address perception distortion in finite block lengths, diffusion is integrated into DJSCC in [22]. Moreover, some works use degraded signals based on diffusion models for semantic communication. A hybrid joint source-channel coding (JSCC) scheme is proposed in [23], where conventional digital communication with the compressed image is complemented with a generative refinement component to enhance the perceptual quality of reconstruction. Invertible Neural Networks are proposed in [24], where the signal is decomposed at the transmitter and is estimated from the degraded part using the diffusion models to recover high-quality source images under extreme conditions at the receiver. The diffusion models are used to denoise in wireless communication. However, current diffusion-based approaches have not fully harnessed the generative capabilities of pre-trained diffusion models [18], which were extensively trained on the LAION-5B dataset [25] with billions of images that required immense GPU resources. Although integrating these pre-trained diffusion models helps preserve their powerful generative capabilities, it presents challenges in flexibly adapting to varying bandwidth constraints and limits their adaptability.
@@
In this paper, we propose a method where the received signal is transformed into the Gaussian-distributed features through forward process in diffusion models, with varying SNRs mapped to different diffusion steps $T$. The received signal is then processed through $T$ iterations of reverse denoising through the Stable Diffusion model to mitigate channel noise. To reduce transmission bandwidth while leveraging the remarkable generative capabilities of pre-trained diffusion models, we integrate a downsampling module for compression and a VAE-based upsampling module at the receiver for reconstruction. The VAE-based module ensures the Gaussian conditions required by the diffusion model and incorporates SNR as a conditioning input to address the sensitivity of feature variance to channel noise. Our contributions can be summarized as follows:
@@
    We propose a communication-efficient generative semantic communication system, which incorporates the forward process of diffusion and channel noise, mapping the channel noise into the $T$-th forward processes in diffusion, adaptable to different signal to noise ratios (SNRs) of the wireless channel. At the receiver, we effectively utilize the diffusion model of the reverse process to remove the channel noise. Multiple kinds of wireless channels, including AWGN, Rayleigh, and multiple-input multiple-output (MIMO) channels, are considered. For the Rayleigh channel, we combine our approach with mean squared error (MSE) equalization, while singular value decomposition (SVD) is utilized for the MIMO channel.
@@
    To reduce bandwidth requirements, we integrate a downsampling module to compress semantic features and a VAE-based upsampling network with reparameterization at the receiver, serving as a plug-and-play module in diffusion-based networks. This approach ensures that the reconstructed features preserve the Gaussian distribution. The VAE-based upsampling module is further designed by incorporating the SNR as a conditioning input for more accurate variance estimation.
@@ III-B Adaptive Forward Process
According to Eqn. (15), when $|\frac{\sqrt{\bar{\alpha}_{u}}\sigma}{\sqrt{1-\bar{\alpha}_{u}}}|=1$, i.e., $\bar{\alpha}_{u}=\frac{1}{\left(1+\sigma^{2}\right)}$, the transmission process over the channel is equivalent to the $u$-th step forward process. This can be expressed as:
$\left.p_{c}\left(\hat{\boldsymbol{s}}\mid\boldsymbol{y}\right)\sim p\left(\boldsymbol{y}_{u}\mid\boldsymbol{y}_{0}\right)\right|_{\bar{\alpha}_{u}=\frac{1}{\left(1+\sigma^{2}\right)}}.$ (16)
@@
$\left.p_{d}\left(\hat{\boldsymbol{s}}_{u-1}\mid\hat{\boldsymbol{s}}_{u},\boldsymbol{\phi}\right)\sim p_{\boldsymbol{\phi}}\left(\boldsymbol{y}_{u-1}\mid\boldsymbol{y}_{u}\right)\right|_{\bar{\alpha}_{u}=\frac{1}{\left(1+\sigma^{2}\right)}},$ (17)
@@
III-C Transformation over Rayleigh and MIMO Channels
To further evaluate the robustness and adaptability of our proposed framework, we extend the transformation of received signals into appropriate features for the forward process of diffusion models to more complex wireless channels, including Rayleigh and MIMO channels.
For the Rayleigh fading channel, MSE equalization is used to transform the received signals into Gaussian-distributed features suitable for the diffusion process. The received signal before equalization is given as:
$\hat{\boldsymbol{s}}=\boldsymbol{h}\boldsymbol{y}+\sigma\cdot\boldsymbol{\epsilon},$ (18)
where $\boldsymbol{h}$ is the channel gain matrix. Using MSE equalization [20], the equalized received signal $\boldsymbol{s}_{\text{MSE}}$ is expressed as:
$\boldsymbol{s}_{\text{MSE}}=\boldsymbol{w}_{\text{MSE}}\hat{\boldsymbol{s}}.$ (19)
The MSE equalization coefficient $\boldsymbol{w}_{\text{MSE}}$ is defined as:
$\boldsymbol{w}_{\text{MSE}}=\frac{\boldsymbol{h}^{*}}{|\boldsymbol{h}|^{2}+\frac{\sigma^{2}}{|\boldsymbol{y}|^{2}}},$ (20)
where $\boldsymbol{h}^{*}$ is the conjugate of $\boldsymbol{h}$. Thus, the received signal after equalization is expressed as:
$\boldsymbol{s}_{\text{MSE}}$ $=\frac{|\boldsymbol{h}|^{2}}{|\boldsymbol{h}|^{2}+\frac{\sigma^{2}}{|\boldsymbol{y}|^{2}}}\cdot\boldsymbol{y}+\sigma\cdot\frac{\boldsymbol{h}^{*}}{|\boldsymbol{h}|^{2}+\frac{\sigma^{2}}{|\boldsymbol{y}|^{2}}}\boldsymbol{\epsilon}$ (21)
$=\frac{|\boldsymbol{h}|^{2}}{|\boldsymbol{h}|^{2}+\frac{\sigma^{2}}{|\boldsymbol{y}|^{2}}}\cdot\left(\boldsymbol{y}+\sigma\cdot\boldsymbol{h}\boldsymbol{\epsilon}\right)$
$=\frac{|\boldsymbol{h}|^{2}}{\sqrt{\bar{\alpha}_{u}}\left(|\boldsymbol{h}|^{2}+\frac{\sigma^{2}}{|\boldsymbol{y}|^{2}}\right)}\cdot\left(\sqrt{\bar{\alpha}_{u}}\cdot\boldsymbol{y}_{0}\right.$
$~{}~{}~{}~{}~{}~{}~{}\left.+\sqrt{1-\bar{\alpha}_{u}}\cdot\frac{\sqrt{\bar{\alpha}_{u}}\sigma\cdot\boldsymbol{h}}{\sqrt{1-\bar{\alpha}_{u}}}\boldsymbol{\epsilon}\right).$
Based on the Eqn. (21), when $|\frac{\sqrt{\bar{\alpha}_{u}}\sigma\cdot\boldsymbol{h}}{\sqrt{1-\bar{\alpha}_{u}}}|=1$, i.e., $\bar{\alpha}_{u}=\frac{1}{1+|\sigma\cdot\boldsymbol{h}|^{2}}$, the transmission process over the Rayleigh channel is equivalent to the $u$-th step forward process in the diffusion process. This equivalence can be expressed as:
$\left.p_{c}\left(\boldsymbol{s}_{\text{MSE}}\mid\boldsymbol{y}\right)\sim p\left(\boldsymbol{y}_{u}\mid\boldsymbol{y}_{0}\right)\right|_{\bar{\alpha}_{u}=\frac{1}{1+|\sigma\cdot\boldsymbol{h}|^{2}}}.$ (22)
For the MIMO channel, we utilize SVD decomposition [36] to decouple the multi-antenna signals into independent streams, ensuring that the transformed signals approximate the Gaussian-distributed features required by the diffusion model. Given the channel state information (CSI) [37, 38], the MIMO channel matrix $\mathbf{H}\in\mathbb{C}^{M\times M}$ is decomposed as follows:
$\boldsymbol{H}=\boldsymbol{U}\boldsymbol{\Sigma}_{\boldsymbol{H}}\boldsymbol{V}^{H},$ (23)
where $\mathbf{U}\in\mathbb{C}^{M\times M}$ is a unitary matrix that satisfies $\mathbf{U}\mathbf{U}^{H}=\mathbf{I}$, representing the left singular vectors of $\mathbf{H}$. Similarly, $\mathbf{V}\in\mathbb{C}^{M\times M}$ is also a unitary matrix, satisfying $\mathbf{V}\mathbf{V}^{H}=\mathbf{I}$, and it represents the right singular vectors of $\mathbf{H}$. $\mathbf{V}^{H}$ is the Hermitian transpose of $\mathbf{V}$. $\boldsymbol{\Sigma}_{\boldsymbol{H}}\in\mathbb{R}^{M\times M}$ is a diagonal matrix, whose entries are the singular values of $\mathbf{H}$, $\sigma_{1},\sigma_{2},\ldots,\sigma_{M}$, arranged in descending order. This is expressed as:
$\boldsymbol{\Sigma}_{\boldsymbol{H}}=\operatorname{diag}\left(\sigma_{1},\sigma_{2},\ldots,\sigma_{M}\right),\quad\sigma_{1}\geq\sigma_{2}\geq\cdots\geq\sigma_{M}\geq 0.$ (24)
The transmitted signal $\boldsymbol{x}_{\text{MIMO}}$ over MIMO channel can be represented as:
$\boldsymbol{x}_{\text{MIMO}}=\mathbf{V}\boldsymbol{y}.$ (25)
The received signal $\boldsymbol{y}_{\text{MIMO}}$ is given as:
$\boldsymbol{y}_{\text{MIMO}}=\boldsymbol{H}\boldsymbol{x}_{\text{MIMO}}+\sigma\cdot\boldsymbol{\epsilon}.$ (26)
By applying the unitary matrix $\mathbf{U}^{H}$ to the received signal $\boldsymbol{y}_{\text{MIMO}}$, the signal is transformed as:
$\boldsymbol{y}_{\text{MIMO}}^{\prime}$ $=\mathbf{U}^{H}\boldsymbol{y}_{\text{MIMO}}=\mathbf{U}^{H}\left(\mathbf{U}\boldsymbol{\Sigma}_{\boldsymbol{H}}\mathbf{V}^{H}\mathbf{V}\boldsymbol{y}+\sigma\cdot\boldsymbol{\epsilon}\right)$ (27)
$=\boldsymbol{\Sigma}_{\boldsymbol{H}}\boldsymbol{y}+\sigma\cdot\boldsymbol{\epsilon}^{\prime},$
where $\boldsymbol{\epsilon}^{\prime}=\mathbf{U}^{H}\boldsymbol{\epsilon}$ is the transformed noise, which remains Gaussian due to the unitary property of $\mathbf{U}$. This separates the $M$-dimensional MIMO channel into $M$ independent scalar sub-channels, represented as:
$\boldsymbol{y}_{i}^{\prime}$ $=\sigma_{i}\cdot\boldsymbol{y}_{i}+\sigma\cdot\boldsymbol{\epsilon}_{i}^{\prime}$ (28)
$=\frac{\sigma_{i}}{\sqrt{\bar{\alpha}_{u}}}\cdot\left(\sqrt{\bar{\alpha}_{u}}\cdot\boldsymbol{y}_{i}+\frac{\sqrt{\bar{\alpha}_{u}}\sigma}{\sqrt{1-\bar{\alpha}_{u}}\sigma_{i}}\sqrt{1-\bar{\alpha}_{u}}\cdot\boldsymbol{\epsilon}_{i}^{\prime}\right),$
where $\sigma_{i}$ is the gain of the $i$-th sub-channel, $i=1,2,\ldots,M$. $\boldsymbol{y}_{i}^{\prime}$ represents the output of the $i$-th sub-channel after transformation. $\boldsymbol{y}_{i}$ represents the transmitted signal over the $i$-th sub-channel. $\boldsymbol{\epsilon}_{i}^{\prime}$ represents the noise component for the $i$-th sub-channel. When $|\frac{\sqrt{\bar{\alpha}_{u}}\sigma}{\sqrt{1-\bar{\alpha}_{u}}\sigma_{i}}|=1$, i.e., $\bar{\alpha}_{u}=\frac{1}{1+\left(\sigma/\sigma_{i}\right)^{2}}$, the transmission process over the sub-channel of MIMO is equivalent to the $u$-th step forward process in the diffusion process. This equivalence can be expressed as:
$\left.p_{c}\left(\boldsymbol{y}_{i}^{\prime}\mid\boldsymbol{y}_{i}\right)\sim p\left(\boldsymbol{y}_{u}\mid\boldsymbol{y}_{0}\right)\right|_{\bar{\alpha}_{u}=\frac{1}{1+\left(\sigma/\sigma_{i}\right)^{2}}}.$ (29)
@@ IV The Details of the Diffusion-Driven Semantic Communication System with VAE-Based Compression
In this section, we present a diffusion-driven semantic communication framework enhanced by VAE-based compression. To ensure the Gaussian distribution of features, we employ VAE-based reparameterization for feature reconstruction. Additionally, we introduce guidance to focus on training the paired upsampling and downsampling networks to reduce training costs. The pre-trained diffusion model integrated with compression leverages guidance from the distribution of the pre-trained model, enabling the efficient generation of corresponding images with reduced bandwidth. Fig. 3 illustrates the framework for detailed semantic communication with compression and guidance.
@@ IV-C The Hybrid Loss from Reparameterization and Guidance
$\mathcal{L}_{v}=\underbrace{\lambda D_{\text{KL}}\left(q_{\boldsymbol{\psi},\boldsymbol{\omega}}\left(\hat{\boldsymbol{y}}\mid\!\boldsymbol{y}\right)\|p_{\boldsymbol{\phi}}\left(\hat{\boldsymbol{y}}\right)\right)}_{\mathcal{L}_{\text{KL}}}+\underbrace{\mathbb{E}\!\left[\left({\boldsymbol{y}^{\prime}\!-\!\boldsymbol{y}}\right)^{2}\!\right]}_{\mathcal{L}_{\text{MSE}}},$ (34)
@@
Proof: The proof is along the lines of the proof in Appendix A and is therefore omitted.
@@ V Results
In this section, we provide a detailed description of the experimental setup and present experimental results to validate the effectiveness of the proposed semantic communication generation with compression and guidance.
V-A Experiment Setup
V-A1 Dataset and Simulation Setting
Our method was trained on the LSUN-Churches dataset, which contains 126000 images, and evaluated on both the LSUN-Churches and LSUN-Bedrooms datasets [40] with 300 images. The pre-trained diffusion model used in our approach is Stable Diffusion v2 [18]. We configure the number of diffusion timesteps to $T=1000$ with a linear schedule, set the learning rate to $1\times 10^{-4}$, and use the training batch size of 4. For the compression module and the VAE-based upsampling module, to reduce the added computational complexity, the residual block is a bottleneck structure composed of three units, which includes a convolutional filter followed by the LeakyReLU and batch normalization. The latent variables of the VAE are modeled to follow a Gaussian distribution through the reparameterization trick. To evaluate the effectiveness of our approach, we transform received signals into appropriate features suitable for the forward process of the diffusion model and extend the transformation to more complex wireless channels, including Rayleigh and MIMO channels. Specifically, we combine our approach with MSE equalization under the assumption that the channel gain $|\boldsymbol{h}|=1$ for the Rayleigh channel, while utilizing SVD decomposition [36] for the MIMO channel with $M=2$. In addition, we also visualized the semantic distinctions between the generated images and real images. This helps to provide a more intuitive understanding of the effectiveness of generating semantic communication based diffusion. The simulations are conducted on the computer equipped with an Intel Xeon Silver 4110 CPU @ 2.10GHz and an NVIDIA RTX A40 GPU.
V-A2 The Metrics
@@ V-B Performance of Channel Integration and Adaptation
TABLE III: The adaptive generation in LSUN-Bedrooms datasets for different channels.
Metric Channel 0dB 3dB 6dB 9dB 12dB
PSNR $\uparrow$ AWGN 22.74 25.59 26.79 26.79 27.83
MIMO 22.74 24.24 25.59 26.79 27.83
Rayleigh 14.87 17.70 21.23 23.27 24.39
SSIM $\uparrow$ AWGN 0.68 0.73 0.76 0.79 0.81
MIMO 0.68 0.73 0.76 0.79 0.81
Rayleigh 0.52 0.57 0.66 0.70 0.72
CLIP $\uparrow$ AWGN 0.88 0.90 0.91 0.91 0.91
MIMO 0.88 0.89 0.90 0.91 0.91
Rayleigh 0.53 0.66 0.82 0.83 0.85
FID $\downarrow$ AWGN 28.43 23.15 19.45 16.56 14.45
MIMO 28.58 23.41 19.38 16.52 14.31
Rayleigh 234.08 115.28 43.26 30.67 22.47
LPIPS $\downarrow$ AWGN 0.22 0.18 0.15 0.12 0.11
MIMO 0.22 0.18 0.15 0.12 0.11
Rayleigh 0.78 0.58 0.34 0.25 0.21
TABLE IV: The adaptive generation in LSUN-Churches datasets for different channels.
Metric Channel 0dB 3dB 6dB 9dB 12dB
PSNR $\uparrow$ AWGN 20.16 21.39 22.52 23.52 24.38
MIMO 20.16 21.39 22.52 23.52 24.38
Rayleigh 13.84 16.15 19.39 21.14 22.01
SSIM $\uparrow$ AWGN 0.58 0.62 0.67 0.70 0.73
MIMO 0.58 0.62 0.67 0.70 0.73
Rayleigh 0.41 0.45 0.55 0.60 0.63
CLIP $\uparrow$ AWGN 0.86 0.88 0.89 0.89 0.89
MIMO 0.86 0.88 0.89 0.89 0.89
Rayleigh 0.45 0.56 0.77 0.82 0.83
FID $\downarrow$ AWGN 27.17 22.43 19.12 16.52 14.72
MIMO 27.10 22.59 19.05 16.63 14.77
Rayleigh 272.50 151.58 43.80 37.96 29.50
LPIPS $\downarrow$ AWGN 0.24 0.20 0.17 0.14 0.13
MIMO 0.24 0.20 0.17 0.14 0.12
Rayleigh 0.86 0.70 0.42 0.28 0.23
Figure 5: The visual results of the adaptation generation across various SNR levels. When the SNR is low, the semantic information of the image can be correctly transmitted, and as the SNR increases, more details can be recovered. Noticeably different regions, highlighted with red boxes, are used for clear comparison to demonstrate how improved SNR contributes to more detailed reconstruction.
@@
We also evaluate the generation performance on the LSUN Bedrooms and LSUN Churches datasets, where the noise from the wireless channel is treated as a complete forward process to adapt to different SNR conditions. To further validate our approach, we extend the transformation to more complex wireless channels, such as Rayleigh and MIMO channels. The received signals are converted into appropriate features suitable for the forward process of the diffusion model. Specifically, we combine our approach with MSE equalization for the Rayleigh channel and utilize SVD decomposition for the MIMO channel. As shown in Tables III and IV, the results demonstrate that the performance of the MIMO channel, after applying SVD decomposition, closely approximates that of the AWGN channel. At high SNR levels, the Rayleigh channel, when combined with MSE equalization, achieves performance comparable to the AWGN channel. These results further validate the adaptability and robustness of our approach across diverse wireless channels. The CLIP-score metrics, used for evaluating the semantic content of the images, demonstrate indicate that diffusion-driven semantic communication ensures accurate transmission of semantic information across various SNR levels. However, at higher SNR, diffusion-driven semantic communication not only ensures accurate transmission of semantic information but also better preserves the details of the images. This result is reflected in the pixel-level evaluation metrics such as PSNR and SSIM. The visual results of generated images at different SNR levels are illustrated in Fig. 5. Even at lower SNR, semantic information can be effectively transmitted. For instance, in the figure, when comparing SNR = 0 dB with SNR = 12 dB, differences (highlighted with red boxes) can be observed in the details of the church’s roof, windows, and architectural style. At lower SNR, the overall semantic content remains consistent, but details are lost. As the SNR increases, these details are better recovered in images.
V-C Performance of Compression with Different Channels and Bandwidths
(a) Comparison of PSNR metric.
(b) Comparison of SSIM metric.
(c) Comparison of LIPIPS metric.
(d) Comparison of CLIP metric.
Figure 6: The comparisons on LSUN-Churches dataset across different wireless channels.
(a) Comparison of PSNR metric.
(b) Comparison of SSIM metric.
(c) Comparison of LIPIPS metric.
(d) Comparison of CLIP metric.
Figure 7: The comparisons on LSUN-Churches dataset across different compression rates.